3.5 Show that the operator*p is not Hermitian (where k and / are positive integers), but that the combination (xd + pl#9/2 is Hermitian.

I've been told properties of commutators are the method of solving this, I'm just unsure of the implementation.

Properties such as: [AB, C] = A[B, C] + [A,C]B and    [A, B] = AB - BA   and   [x, p] = i * h(bar)    and   [p, x2 ] = -2i(hbar)x
Answer #1

The property we can use here is that

[A,B^n]=n[A,B]B^{n-1} and[x,p_x]=i\hbar and [p_x,x]=-i\hbar

[x^k,p_x^l]=kx^{k-1}[x,p_x^l]=kx^{k-1}[x,p_x]lp_x^{l-1}=i\hbar klx^{k-1}p_x^{l-1}\ne 0.................(1)

Since it is not equal to zero so it is not Hermitian.

Now solving for the other combination.

\frac{1}{2}[x^kp_x^l+p_x^lx^k]=\frac{1}{2}[x^kp_x^l]+\frac{1}{2}[p_x^l,x^k]

Now we can use the same formulation as above. The first term is the same as equation (1) multiplied by 1/2. And in second term x and px are interchanged. In that case, using the property mention above we will come out with a minus sign. Therefore

\Rightarrow i\hbar klx^{k-1}p_x^{l-1}-i\hbar klp_x^{l-1}x^{k-1}=0

Therefore it is Hermitian.